OpenStudy (anonymous):

Input the equation of the given line. The line including (2, 3) and (2, 5).

hartnn (hartnn):

could you find slope ?

OpenStudy (anonymous):

5-3=2 2-2=0 slope=2/0

hartnn (hartnn):

yes! infinite slope implies this is a vertical line!

OpenStudy (austinl):

Ok, the line will have a slope of, \(\displaystyle \text{slope}=\frac{\text{rise}}{\text{run}}=\frac{y_2-y_1}{x_2-x_1}\) in the form, \(y=mx+b\) That will give you the slope. Then you can take the \((x,y)\rightarrow(2,3)\) points given and solve for the y-intercept \(b\). Make sense?

OpenStudy (austinl):

And I typed all of that out and @hartnn already swooped in lol

hartnn (hartnn):

and equation of vertical line is of the form x =c where 'c' is the 'x' co-ordinate lol

hartnn (hartnn):

what is 'x' co-ordinate here ?

OpenStudy (anonymous):

2 and 2

hartnn (hartnn):

so the equation is just x=2 ! thats it :)

hartnn (hartnn):

doubts ?